40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Feb 13 16:35:32 2020, anchal, DailyProgress, BEAT, BN Detector was saturated. Reduced laser powers. 
    Reply  Fri Feb 14 10:59:13 2020, anchal, DailyProgress, BEAT, Beatnote noise lowered in low frequency region 
       Reply  Mon Feb 17 16:08:23 2020, anchal, DailyProgress, BEAT, Beatnote noise lowered in low frequency region with ISS CTN_Noise_Budget_Effect_ISS_ON.pdf
          Reply  Fri Feb 21 16:22:04 2020, anchal, DailyProgress, BEAT, Beatnote noise lowered in low frequency region with ISS CTN_Noise_Budget_Effect_ISS_ON_New.pdf
             Reply  Tue Mar 3 10:02:12 2020, anchal, DailyProgress, BEAT, Lowest ever beatnote spectrum today! CTN_Noise_Budget_March03-2020.pdf
                Reply  Thu Mar 12 11:06:35 2020, anchal, DailyProgress, BEAT, Lowest ever beatnote spectrum today! CTN_Noise_Budget_March11-2020.pdf
                   Reply  Wed Mar 18 12:51:22 2020, anchal, DailyProgress, BEAT, Super beatnote measurement analysis results CTN_Beatnote_SuperMeasurement_March13-16.pdf
                      Reply  Mon Apr 13 16:52:30 2020, anchal, DailyProgress, BEAT, Beatnote measurements back on track CTN_Latest_BN_Spec.pdf
Message ID: 2535     Entry time: Fri Feb 14 10:59:13 2020     In reply to: 2532     Reply to this: 2540
Author: anchal 
Type: DailyProgress 
Category: BEAT 
Subject: Beatnote noise lowered in low frequency region 

With some of the changes done recently, the beatnote noise has lowered in the low-frequency region indicating reduction in scatter.


Latest BN Spectrum: CTN_Latest_BN_Spec.pdf

Daily BN Spectrum: CTN_Daily_BN_Spec.pdf


Relevant elog post:

CTN:2530 : Increased sampling rate to 125kSa/s; lowest noise in higher frequencies

CTN:2533 : blocked NF1811 with hex beam dump

CTN:2535 : BN Detector was saturated. Reduced laser powers.

CTN:2531 : Further iterated back and forth to optimized FSS Gains.

ELOG V3.1.3-